Cassandra has a loan of $35,720. This loan has a simple interest rate of 4% per year. What is the amount of interest Cassandra will pay on this loan at the end of two years?
Answer by math_tutor2020(3817)


Simple Interest Formula
i = P*r*t

We have:
i = unknown interest
P = 35720 = loan amount (you can think of it like a deposit)
r = 0.04 = decimal form of annual interest rate
t = 2 = number of years

We then compute the following:
i = P*r*t
i = 35720*0.04*2
i = 2857.60

Answer: $2,857.60